OpenStudy (anonymous):

hoose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (anonymous):

the one that looks like \[y+1=4(x+3)\]

OpenStudy (anonymous):

ok can you show me how

OpenStudy (anonymous):

sure

OpenStudy (jackellyn):

@boycrazzy you can use this formula y – y1 = m(x – x1) using the points given to you in the problem (x1,y1) and the slope, m.

OpenStudy (anonymous):

if your choices look like \(y=mx+b\) then we can change that form easily enough multiply out on the right \[y+1=4(x+3)\] to get \[y+1=4x+12\] then subtract 1 from both sides and get \[y=4x+11\]

OpenStudy (anonymous):

ok i get that but it was -1 and -3

OpenStudy (anonymous):

if you are asking where the equation comes from, you have a point \((-3,-1)\) and a slope \(m=4\) make replacements in the point slope formula \[y-y_1=m(x-x_1)\] to get \[y-(-1)=4(x-(-3))\] which you would really write as \[y+1=4(x+3)\]

OpenStudy (anonymous):

oh lol that makes perfect sense just wasn't think thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

don't forget the minus sign is in the formula if one of the numbers happens to be negative, there is then going to be two minus signs